Determine zα for the following. (Round your answers to two decimal places.)

(a)    α = 0.0089

(b)    α = 0.09

(c)    α = 0.707

Please explain where and how you got the answer.

Thank you.
Answer #1
Concepts and reason

The concept of standard normal variate is used to solve this problem.

The standard normal variate is a normal distribution which has a mean equal to zero and a standard deviation equal to one, and the curve of this variate is bell-shaped.

Fundamentals

The confidence level and the level of significance are related with the following relationship:

Confidencelevel=1levelofsignificance{\rm{Confidence level}} = {\rm{ }}1 - {\rm{ level of significance}}

The confidence level shows the area to the left of the corresponding z-value, whereas the level of significance shows the area to the right of the corresponding z-value.

(a)

To find the value of Zα{Z_\alpha } at the significance level 0.0089, the confidence level would be;

10.0089=0.99111 - 0.0089 = 0.9911

Then use Excel function normsinv(0.9911){\rm{normsinv}}\left( {0.9911} \right) to get the corresponding z- value. The Excel formula used is as shown:

1 =NORMSINV(0.9911)
2

(b)

To find the value of Zα{Z_\alpha } at the significance level 0.09, the confidence level would be;

10.09=0.911 - 0.09 = 0.91

Use the Excel function normsinv(0.91){\rm{normsinv}}\left( {0.91} \right) to get the Z value as shown:

1 =NORMSINV(0.9911)
2
3 ENORMSINV(0.91)

(c)

To find the value of Zα{Z_\alpha } at the significance level 0.707, the confidence level would be:

10.707=0.2931 - 0.707 = 0.293

Use the Excel function normsinv(0.293){\rm{normsinv}}\left( {0.293} \right) to get the Z value as shown:

1 ENORMSINV(0.9911)
2.
3 =NORMSINV(0.91)
4
5 =NORMSINV(0.293)
6

Ans: Part a

The Z-value for a significance level of 0.0089 is 2.372.37 .

Part b

The Z-value for a significance level of 0.91 is 1.341.34 .

Part c

The Z-value for a significance level of 0.293 is 0.545 - 0.545 .